# SoK: Speedy Secure Finality

### v1

@@ Abstract
While Ethereum has successfully achieved dynamic availability along with safety, a fundamental delay remains between transaction execution and immutable finality. In Ethereum’s current Gasper protocol, this latency extends to approximately 15 minutes, exposing the network to ex-ante reorg attacks, leading to MEV capture and limiting the efficiency of economic settlement. This inherent delay along with cracks in Gasper has catalyzed the emergence of ”Speedy Secure Finality” (SSF) research—efforts dedicated to minimizing confirmation times without compromising rigorous security guarantees.
@@
This work presents a survey of the state of the art on the evolution of fast finality protocols. We define the core theoretical primitives necessary for understanding this landscape, including reorg resilience and the generalized sleepy model, and trace their development from Goldfish to RLMD-GHOST. We then analyze the practical trade-offs of Single Slot Finality against the communication bottlenecks of large validator sets. Finally, we survey the 3-Slot Finality (3SF) protocol as a pragmatic convergence of these efforts, bridging the gap between theoretical safety guarantees and the engineering constraints of the Ethereum network.
1 Introduction
@@
Consensus protocols enable a distributed system vis-à-vis blockchains to maintain a consistent replicated state in the presence of faults, adversarial behavior, and network uncertainty. In blockchain systems, consensus further serves as the mechanism by which a globally ordered ledger is constructed under open participation and economic incentives. A central property governing the security and usability of such systems is finality: the guarantee that once a block is confirmed, it will not be reverted in any admissible execution. A detailed discussion about consensus protocols can be found in [bano2019sok, arafatsok].
@@
Over the past two decades, consensus protocols have evolved along multiple, often conflicting design axes. Classical Byzantine Fault Tolerant (BFT) (see [zhong2023byzantine] for a comprehensive survey) protocols achieve deterministic finality under partial synchrony by assuming a fixed replica set with a bounded fraction of Byzantine faults. Nakamoto style longest chain protocols, by contrast, operate under dynamic participation and unknown resource levels, achieving adaptivity and liveness at the cost of replacing finality with probabilistic confirmation. This tension is fundamental where Lewis-Pye and Roughgarden [resourcepool] formalize an impossibility result showing that adaptivity and finality cannot be simultaneously achieved in the unsized, partially synchronous setting.
@@
Modern proof of stake blockchains (see [saleh2021blockchain]) combine elements from both paradigms that separate availability from finality. Ethereum’s consensus protocol, Gasper [gasper], is the most prominent example, pairing a dynamically available fork-choice rule with a BFT-style finality gadget in an ebb-and-flow architecture. While such designs have proven robust in practice, they introduce choices between confirmation latency, reorganization risk, and implementation complexity. Understanding these trade-offs, and the extent to which they can be optimized is the focus of this paper.
At a foundational level, consensus in blockchains can be understood through the lens of state machine replication (SMR), a classical abstraction in distributed systems (see related work [marandi2011high, distler2021byzantine]). In SMR, a deterministic state machine $M=(S,\Sigma,\delta)$ is replicated across multiple processes, and the goal is to ensure that all correct replicas apply the same totally ordered sequence of inputs. Agreement on this order suffices to guarantee that replicas remain in identical states, even in the presence of faults.
Classical SMR protocols, such as Practical Byzantine Fault Tolerance (pBFT), achieve agreement by enforcing total order broadcast under partial synchrony. They assume a fixed replica set and guarantee safety through quorum intersection: once a value is committed by a quorum, it cannot be reverted, yielding immediate finality. However, these guarantees rely critically on static participation and continuous availability.
Permissionless blockchains depart from this setting by allowing open, resource-weighted participation and dynamic validator availability. These features invalidate the static assumptions underlying classical SMR protocols and motivate alternative consensus models, including longest chain protocols and sleepy-consensus variants, which relax participation assumptions while modifying or weakening finality guarantees.
@@
Fast finality (see [anceaume2020finality] and applications in [chou2025minimmit] and [pan2021plume]) refers to protocols that aim to minimize the latency between block proposal and irrevocable confirmation, subject to the constraints imposed by participation dynamics and network assumptions. Ethereum’s Gasper protocol achieves finality on the order of tens of slots, creating a window during which blocks are vulnerable to reorganization and economically motivated attacks. Recent work seeks to shrink this window by refining fork-choice rules, vote handling, and participation models, without abandoning dynamic availability. As far as the authors know, there is no concrete documentation of the fast finality design model in blockchains and this SoK systematically surveys this space through the taxonomy outlined above.
@@
We begin by formalizing the relevant network and fault models, including partial synchrony and the sleepy model. We then analyze Ethereum’s Gasper protocol as a canonical ebb-and-flow construction. Building on this foundation, we examine recent proposals for faster finality, including Goldfish, RLMD-GHOST, and single-slot and three-slot finality protocols. For each class of designs, we identify the underlying assumptions, formal guarantees, and implementation constraints that govern their viability in large-scale permissionless systems.
@@ 2.1 Network Assumptions
Conversely, if a node is adversarial, it may wish to withhold that message for some definite or indefinite amount of time for malicious purposes. Both of these scenarios are indistinguishable and hence, it is necessary to assume some message delay to formalize any consensus protocol. Based on these assumption, a network can be classified into three models.
Definition 2.1 (Synchronous Network).
@@
Definition 2.2 (Asynchronous Network).
@@
As a result of this assumption, these protocols are often very robust but are more complex and harder to reason about. Such models are subject to fundamental constraints, most notably the FLP impossibility result. [flpimposibility]
Definition 2.3 (Partially-Synchronous Network).
A partially-synchronous network is one where $\Delta$ is a finite and known time bound following a period known as the global stabilization time GST. [gst]
@@
2.2 Practical Byzantine Fault Tolerance (pBFT)
Practical Byzantine Fault Tolerance (pBFT) was introduced by Castro and Liskov [pbft] in 1999 to achieve fault-tolerant state machine replication in an adversarial environment. It replicates a deterministic state machine across $n$ replicas of which at most $f$ may be Byzantine. The network is partially synchronous but supports authenticated reliable channels. The protocol’s fault tolerance limit is derived from the necessity to ensure a majority quorum of honest nodes under the worst-case scenario. Assuming $f$ nodes exhibit Byzantine behavior, up to $f$ distinct nodes may incur non-adversarial crash faults. The following summarizes the core mechanics of the protocol as defined in [pbft].
Theorem 2.1 (Fault bound).
@@ 2.2 Practical Byzantine Fault Tolerance (pBFT)
The protocol operates by forming quorums of size $2f+1$. If $Q_{1}$ and $Q_{2}$ are such quorums, then $|Q_{1}\cap Q_{2}|\geq f+1$ whenever $n\geq 3f+1$. Since at most $f$ replicas are faulty, every quorum intersection contains at least one honest replica. This property is necessary for agreement and becomes impossible when $n\leq 3f$; hence the bound is tight. ∎
@@ 2.2.3 Commit
Definition 2.5 (Committed-Local).
@@
Theorem 2.2 (Safety).
@@
Together, safety and liveness establish pBFT as a complete Byzantine fault-tolerant replication protocol under the $n\geq 3f+1$ bound and partial synchrony. Its reliance on a fixed, fully active replica set stands in contrast to consensus models with dynamic or intermittently online participants. We next examine the Sleepy Model of Consensus, which relaxes these assumptions and enables agreement in settings with fluctuating node availability.
2.3 Sleepy Model of Consensus
In permissionless blockchain settings, replicas need not remain continuously active. Pass and Shi [sleepy] formalize this by introducing the notion of sleepy behavior, in contrast to the fully synchronous participation assumed by classical BFT protocols such as pBFT.
Definition 2.6 (Sleepy Node [sleepy]).
@@ 2.3 Sleepy Model of Consensus
Pass and Shi introduce a consensus model [sleepy] in which the relevant security parameter is the fraction of honest online replicas at each time step. Their protocol achieves safety provided that adversarial online participation never exceeds the honest-online majority and achieves liveness under partial network synchrony given the same condition. This relaxes the static quorum requirement and makes consensus compatible with fluctuating availability. The Ouroboros proof-of-stake protocol [ouroboros] later termed this robustness property dynamic availability.
Definition 2.7 (Dynamic Availability).
@@ 2.4 Finality and Adaptivity
Nakamoto demonstrated in the Bitcoin whitepaper[bitcoin] that the probability that an attacker can re-org a block decreases exponentially as new blocks are confirmed on the canonical chain. This sort of guarantee is called probabilistic finality. We follow the framework of Lewis-Pye and Roughgarden [resourcepool] to formalize the notions of finality and adaptivity for permissionless consensus protocols.
@@
By contrast, longest-chain Nakamoto-style protocols (e.g., Bitcoin) do not satisfy finality: confirmation is only secure under ongoing participation of an honest majority in a sufficiently synchronous regime. As shown in the Bitcoin whitepaper [bitcoin], for an attacker with a resource fraction $q<\tfrac{1}{2}$, the probability that a $k$-deep block can be rolled back decays exponentially in $k$, yielding probabilistic finality but not finality in the above sense.
@@
Protocols in [resourcepool] are analyzed relative to a resource pool $R$, which assigns to each public key $U$ and timeslot $t$ a nonnegative resource balance $R(U,t,M)$ (e.g., hashrate in PoW or stake in PoS), possibly depending on the message state $M$. Let
@@
Theorem 2.4 (Impossibility of adaptivity and finality [resourcepool, Thm. 4.1]).
@@
The intuition is that, in the unsized and partially synchronous setting, an extended period of network asynchrony is indistinguishable (from the protocol’s local perspective) from a sharp drop in the total resource pool. Liveness in the unsized setting forces the protocol to keep confirming blocks even during such periods, while security (and hence finality) forbids confirmation during sufficiently bad partitions. These two requirements are incompatible, yielding Theorem 2.4.
@@
  • Longest-chain Nakamoto-style protocols (Bitcoin, Snow White, etc.) are adaptive in the unsized setting but, by Theorem 2.4, cannot have finality. They only offer probabilistic finality in the sense of exponentially decreasing reorg probability as confirmations accumulate.
@@
2.5 Ebb n Flow Protocols
Since the properties of availability and finality cannot co-exist in a single consensus protocol, Nue, Tas and Tse [ebbnflow] formulated a new class of flexible consensus protocols called ebb-n-flow protocols.
Definition 2.12 (Ebb-n-Flow Protocols).
@@ 2.5 Ebb n Flow Protocols
Ethereum’s consensus protocol Gasper is an ebb-n-flow protocol that uses LMD-GHOST as its availability protocol and Casper as its finality gadget. It considers a partially synchronous network, where honest nodes are sleepy until global awake time GAT, [ebbnflow] after which all nodes are awake. It also assumes that adversaries are always awake. Then
@@
3 The Ethereum Consensus Protocol
Ethereum is the most decentralized blockchain in the world. It uses a Proof of Stake model as its consensus protocol. This model is called Gasper [gasper], an ebb-n-flow protocol which can provide safety, plausible liveness, and probabilistic liveness under different assumptions. Gasper combines Casper (the friendly-finality gadget) [casper] and LMD-GHOST (fork-choice rule).
@@ 3 The Ethereum Consensus Protocol
3.1 LMD-GHOST: The Fork-Choice Rule of Gasper
LMD-GHOST is the fork-choice rule for the Ethereum consensus protocol. It comprises of two acronyms that stand for ”Latest Message Driven” and ”Greedy Heaviest-Observed Sub-Tree”. LMD-GHOST is an adaptation of the GHOST protocol introduced by Sompolinsky and Zohar [ghost].
Definition 3.1 (Fork-Choice Rule).
@@ 3.1 LMD-GHOST: The Fork-Choice Rule of Gasper
As the name suggests, LMD-GHOST is a Greedy algorithm that runs recursively on a tree. Starting at $B_{genesis}$ or any block on which the rule is to be run, we select the heaviest branch descending from it, i.e. we choose the child block from the root which has the highest weight. In a situation where two or more child blocks have the same weight, the block with the greatest block-hash will be chosen. We then repeat the process recursively with the chosen child block as the new root until we reach a leaf. For a given $V$, this leaf block $B_{V}$ is the output of the algorithm and will be considered the head of the chain.
@@
3.2 Casper: The Friendly Finality Gadget
@@
Theorem 3.1 (Accountable Safety).
@@
Theorem 3.2 (Plausible Liveness).
@@
3.3 Gasper: Ethereum’s Ebb-n-Flow Protocol

### v2

@@ Abstract
While Ethereum has successfully achieved dynamic availability along with safety, a fundamental delay remains between block production and immutable finality. In Ethereum’s current Gasper protocol, this latency extends to approximately 15 minutes, exposing the network to ex-ante reorg attacks, leading to MEV capture and limiting the efficiency of economic settlement. This inherent delay along with cracks in Gasper has catalyzed the emergence of “Speedy Secure Finality” (SSF): research efforts dedicated to minimizing confirmation times without compromising rigorous security guarantees. This work presents a survey of the state of the art on the evolution of fast finality protocols. We define the core theoretical primitives necessary for understanding this landscape, including reorg resilience and the generalized sleepy model, and trace their development from Goldfish to RLMD-GHOST. We then analyze the practical trade-offs of Single Slot Finality against the communication bottlenecks of large validator sets. Finally, we survey the 3-Slot Finality (3SF) protocol as a pragmatic convergence of these efforts, bridging the gap between theoretical safety guarantees and the engineering constraints of the Ethereum network.
I Introduction
@@
Consensus protocols enable a distributed system vis-à-vis blockchains to maintain a consistent replicated state in the presence of faults, adversarial behavior, and network uncertainty. In blockchain systems, consensus further serves as the mechanism by which a globally ordered ledger is constructed under open participation and economic incentives. Two central properties governing the security and usability of such systems are availability: the guarantee that block production does not halt as long an honest majority of validators are online, and finality: the guarantee that once a block is confirmed, it will not be reverted in any admissible execution. A detailed discussion about consensus protocols can be found in [3, 2].
@@
Over the past two decades, consensus protocols have evolved along multiple, often conflicting design axes. Classical Byzantine Fault Tolerant (BFT) (see [29] for a comprehensive survey) protocols achieve deterministic finality under partial synchrony by assuming a fixed replica set with a bounded fraction of Byzantine faults. Nakamoto style longest chain protocols, by contrast, operate under dynamic participation and unknown resource levels, achieving adaptivity and liveness at the cost of replacing finality with probabilistic confirmation. This tension is fundamental where Lewis-Pye and Roughgarden [19] formalize an impossibility result showing that adaptivity and finality cannot be simultaneously achieved in the unsized, partially synchronous setting.
@@
Modern proof of stake blockchains [26] combine elements from both paradigms that separate availability from finality. Ethereum’s consensus protocol, Gasper [5], is the most prominent example, pairing a dynamically available fork-choice rule with a BFT-style finality gadget in an ebb-and-flow architecture. While such designs have proven robust in practice, they introduce choices between confirmation latency, reorganization risk, and implementation complexity. Understanding these trade-offs, and the extent to which they can be optimized is the focus of this paper.
@@
Fast finality (see [1] and applications in [9] and [24]) refers to protocols that aim to minimize the latency between block proposal and irrevocable confirmation, subject to the constraints imposed by participation dynamics and network assumptions. Ethereum’s Gasper protocol achieves finality on the order of tens of slots, creating a window during which blocks are vulnerable to reorganization and economically motivated attacks. Recent work seeks to shrink this window by refining fork-choice rules, vote handling, and participation models, without abandoning dynamic availability. As far as the authors know, there is no concrete documentation of the fast finality design model in blockchains and this SoK systematically surveys this space through the taxonomy outlined above.
@@
We begin by formalizing the relevant network and fault models, including partial synchrony and the sleepy model. We then analyze Ethereum’s Gasper protocol as a canonical ebb-and-flow construction. Building on this foundation, we examine recent proposals for faster finality, including Goldfish, RLMD-GHOST, single-slot finality and then three-slot finality protocols. For each class of designs, we identify the underlying assumptions, formal guarantees, and implementation constraints that govern their viability in large-scale permissionless systems.
@@ II-A Network Assumptions
Conversely, if a node is adversarial, it may wish to withhold that message for some definite or indefinite amount of time for malicious purposes. Both of these scenarios are indistinguishable and hence, it is necessary to assume some message delay to formalize any consensus protocol. Based on these assumption, a network can be classified into three models [20]:
Definition II.1 (Synchronous Network).
@@
Definition II.2 (Asynchronous Network).
@@
As a result of this assumption, these protocols are often very robust but are more complex and harder to reason about. Such models are subject to fundamental constraints, most notably the Fischer-Lynch-Patterson (FLP) impossibility result [17].
Definition II.3 (Partially-Synchronous Network).
A partially-synchronous network is one where $\Delta$ is a finite and known time bound following a period known as the global stabilization time GST [16]
@@
II-B Practical Byzantine Fault Tolerance (pBFT)
Practical Byzantine Fault Tolerance (pBFT) was introduced by Castro and Liskov [8] in 1999 to achieve fault-tolerant state machine replication in an adversarial environment. It replicates a deterministic state machine across $n$ replicas of which at most $f$ may be Byzantine. The network is partially synchronous but supports authenticated reliable channels. The protocol’s fault tolerance limit is derived from the necessity to ensure a majority quorum of honest nodes under the worst-case scenario. Assuming $f$ nodes exhibit Byzantine behavior, up to $f$ distinct nodes may incur non-adversarial crash faults. The following summarizes the core mechanics of the protocol as defined in [8].
Theorem II.1 (Fault bound).
@@ II-B Practical Byzantine Fault Tolerance (pBFT)
The protocol operates by forming quorums of size $2f+1$. If $Q_{1}$ and $Q_{2}$ are such quorums, then $|Q_{1}\cap Q_{2}|\geq f+1$ whenever $n\geq 3f+1$. Since at most $f$ replicas are faulty, every quorum intersection contains at least one honest replica. The bound $n\geq 3f+1$ guarantees that for any two quorums $Q_{1},Q_{2}\subseteq[n]$ with $|Q_{1}|=|Q_{2}|=2f+1$, we have $|Q_{1}\cap Q_{2}|\geq f+1$, ensuring the presence of at least one honest replica in every quorum intersection. This property is necessary for agreement and becomes impossible when $n\leq 3f$; hence the bound is tight. ∎
@@ II-B3 Commit
Definition II.5 (Committed-Local).
@@
Figure 1: Normal case execution with Replica 0 as leader [8]
Theorem II.2 (Safety).
@@
Together, safety and liveness establish pBFT as a complete Byzantine fault tolerant replication protocol under the $n\geq 3f+1$ bound and partial synchrony. Its reliance on a fixed, fully active replica set stands in contrast to consensus models with dynamic or intermittently online participants. We next examine the Sleepy Model of Consensus, which relaxes these assumptions and enables agreement in settings with fluctuating node availability.
II-C Sleepy Model of Consensus
In permissionless blockchain settings, replicas need not remain continuously active. Pass and Shi [25] formalize this by introducing the notion of sleepy behavior, in contrast to the fully synchronous participation assumed by classical BFT protocols such as pBFT.
Definition II.6 (Sleepy Node [25]).
@@ II-C Sleepy Model of Consensus
Pass and Shi introduce a consensus model in which the relevant security parameter is the fraction of honest online replicas at each time step. Their protocol achieves safety provided that adversarial online participation never exceeds the honest-online majority and achieves liveness under partial network synchrony given the same condition. This relaxes the static quorum requirement and makes consensus compatible with fluctuating availability. The Ouroboros proof-of-stake protocol [18] later termed this robustness property dynamic availability.
Definition II.7 (Dynamic Availability).
@@ II-D Finality and Adaptivity
Nakamoto demonstrated in the Bitcoin whitepaper[21] that the probability that an attacker can re-org a block decreases exponentially as new blocks are confirmed on the canonical chain. This sort of guarantee is called probabilistic finality. We follow the framework of Lewis-Pye and Roughgarden [19] to formalize the notions of finality and adaptivity for permissionless consensus protocols.
@@
By contrast, longest-chain Nakamoto-style protocols (e.g., Bitcoin) do not satisfy finality: confirmation is only secure under ongoing participation of an honest majority in a sufficiently synchronous regime. As shown in the Bitcoin whitepaper [21], for an attacker with a resource fraction $q<\tfrac{1}{2}$, the probability that a $k$-deep block can be rolled back decays exponentially in $k$, yielding probabilistic finality but not finality in the above sense.
@@
Protocols in [19] are analyzed relative to a resource pool $R$, which assigns to each public key $U$ and timeslot $t$ a nonnegative resource balance $R(U,t,M)$ (e.g., hashrate in PoW or stake in PoS), possibly depending on the message state $M$. Let
@@
Theorem II.4 (Impossibility of adaptivity and finality [19, Thm. 4.1]).
@@
The intuition is that, in the unsized and partially synchronous setting, an extended period of network asynchrony is indistinguishable (from the protocol’s local perspective) from a sharp drop in the total resource pool. Liveness in the unsized setting forces the protocol to keep confirming blocks even during such periods, while security (and hence finality) forbids confirmation during sufficiently bad partitions. These two requirements are incompatible, yielding Theorem II.4.
@@
  • Longest-chain Nakamoto-style protocols (Bitcoin, Snow White, etc.) are adaptive in the unsized setting but, by Theorem II.4, cannot have finality. They only offer probabilistic finality in the sense of exponentially decreasing reorg probability as confirmations accumulate.
@@
II-E Ebb n Flow Protocols
Since the properties of availability and finality cannot co-exist in a single consensus protocol, Nue, Tas and Tse [22] formulated a new class of flexible consensus protocols called ebb-n-flow protocols.
Definition II.12 (Ebb-n-Flow Protocols).
@@ II-E Ebb n Flow Protocols
Figure 2: (a) uses a simple SMR protocol $\Pi$ to output a single ledger LOGLOG (b) uses an ebb-n-flow protocol $\Pi_{sac}$ made up of an availability protocol $\Pi_{lc}$ and a finality protocol $\Pi_{bft}$ which outputs an availability $LOG_{da}$ and finality $LOG_{fin}$ ledger respectively [22]
@@
Ethereum’s consensus protocol Gasper is an ebb-n-flow protocol that uses LMD-GHOST as its availability protocol and Casper as its finality gadget. It considers a partially synchronous network, where honest nodes are sleepy until global awake time GAT, [22] after which all nodes are awake. It also assumes that adversaries are always awake. Then
@@
III The Ethereum Consensus Protocol
Ethereum is the most decentralized blockchain in the world. It uses a Proof of Stake model as its consensus protocol. This model is called Gasper [5], an ebb-n-flow protocol which can provide safety, plausible liveness, and probabilistic liveness under different assumptions. Gasper combines Casper (the friendly-finality gadget) [4] and LMD-GHOST (fork-choice rule).
@@ III The Ethereum Consensus Protocol
III-A LMD-GHOST: The Fork-Choice Rule of Gasper
LMD-GHOST is the fork-choice rule for the Ethereum consensus protocol. It comprises of two acronyms that stand for ”Latest Message Driven” and ”Greedy Heaviest-Observed Sub-Tree”. LMD-GHOST is an adaptation of the GHOST protocol introduced by Sompolinsky and Zohar [28].
Definition III.1 (Fork-Choice Rule).
@@ III-A LMD-GHOST: The Fork-Choice Rule of Gasper
LMD-GHOST is a Greedy algorithm that runs recursively on a tree. Starting at $B_{genesis}$ or any block on which the rule is to be run, we select the heaviest branch descending from it, i.e. we choose the child block from the root which has the highest weight. In a situation where two or more child blocks have the same weight, the block with the greatest block-hash will be chosen. We then repeat the process recursively with the chosen child block as the new root until we reach a leaf. For a given $V$, this leaf block $B_{V}$ is the output of the algorithm and will be considered the head of the chain.
@@
III-B Casper: The Friendly Finality Gadget
@@
Theorem III.1 (Accountable Safety).
@@
Theorem III.2 (Plausible Liveness).
@@
III-C Gasper: Ethereum’s Ebb-n-Flow Protocol